OpenStudy (toxicsugar22):

The density of balsa wood is 7.8 pounds per cubic foot. What is the weight, in kilograms, of a piece of balsa wood 4.0 inches by 6.0 inches by 20.0 inches.

OpenStudy (anonymous):

density = mass/volume do you know the volume of the piece in square inches?

OpenStudy (toxicsugar22):

is it 480

OpenStudy (toxicsugar22):

the volume

OpenStudy (australopithecus):

correct

OpenStudy (australopithecus):

do you know what units it is in?

OpenStudy (australopithecus):

the volume?

OpenStudy (toxicsugar22):

480 c^3

OpenStudy (australopithecus):

in^3

OpenStudy (australopithecus):

inches cubed

OpenStudy (australopithecus):

no idea what the c is for

OpenStudy (australopithecus):

all units work like this btw

OpenStudy (australopithecus):

ok so all you have to do is convert cubic inches to cubic feet

OpenStudy (australopithecus):

because we know density of the wood is 7.8lbs/ft^3

OpenStudy (australopithecus):

we always want to work with like units hen trying to figure stuff like this out

OpenStudy (australopithecus):

can you convert it over or do you need me to explain how?

OpenStudy (toxicsugar22):

expplain please

OpenStudy (australopithecus):

I'm going to show you exactly the same thing I showed you last time hopefully it makes sense to you this time :D 12 inch = 1 feet Convert them to squared units 12 inch*12inch = 1 feet*1feet = 144inch^2 = 1 feet^2 Now convert them to cubic units 144inch^2*12inch = 1 feet^2*1 feet = 1728inch^3 = 1 feet^3 If we wanted to convert them to the power 4 we would keep doing this, 1728inch^3 * 12inch = 1 feet^3*1 feet = 20736inch^4 = 1 feet^4 So in summery to get to cubic units from just regular feet and inches 12 inch*12inch*12inch = 1 feet*1 feet*1 feet = 1728inch^3 = 1 feet^3 Next to convert over we simply make a ratio up we want to eliminate inches and be left with feet so we put feet in the numerator and put inches in the denominator. Think of the units kind of like numbers when you divide something by itself you always get 1 which when multiplied by anything just gives back anything unchanged. For example 1/1 = 1 3333/3333 = 1 43434/43434 = 1 etc. \[\frac{{1feet^3}}{1738inch^3}*480inch^3\] \[\frac{1feet^3*480inch^3}{1738inch^3}\]

OpenStudy (australopithecus):

get it?

OpenStudy (toxicsugar22):

yesh thanks

OpenStudy (toxicsugar22):

so anwer would be .276

OpenStudy (toxicsugar22):

right

OpenStudy (australopithecus):